Find (a) the mean, (b) the median, (c) the mode, and (d) the midrange.



Grade Point Averages. The average undergraduate grade point average (GPA) for the 25


top-ranked medical schools is listed below


3.80 3.77 3.70 3.74 3.70


3.86 3.76 3.68 3.67 3.57


3.83 3.70 3.80 3.74 3.67


3.78 3.74 3.73 3.65 3.66


3.75 3.64 3.78 3.73 3.64

1
Expert's answer
2022-06-06T04:20:39-0400

Sorted Data Set: 


"3.57, 3.64, 3.64, 3.65, 3.66,"

"3.67, 3.67, 3.68, 3.70, 3.70,"

"3.70, 3.73, 3.73, 3.74, 3.74,"

"3.74, 3.75, 3.76, 3.77, 3.78,"

"3.78, 3.80, 3.80, 3.83, 3.86"

(a)


"mean=\dfrac{1}{25}(3.57+ 3.64+3.64+3.65+3.66"




"+ 3.67+3.67+3.68+3.70+3.70+3.70"



"+3.73+3.73+3.74+3.74+3.74+3.75"

"3.76+3.77+3.78+3.78+3.80+3.80"


"+3.83+3.86=3.7236"

(b)



"median=3.73\ (13^{th} \ place)"

(c)



"mode=3.70, 3.74,"

each appeared 3 times


(d)



"midrange=\dfrac{3.57+3.86}{2}=3.715"